The owner of a store sells raisins for $3.20 per pound and nuts for $2.40. He decides to mix the raisins and nuts to sell 50lb. of the mixture for $2.72 per pound. What quantities of raisins and nuts should he use?
Answer by mananth(16946)

Price ---------------- quantity
raisins 3.20 ---------------- x pounds
nuts 2.40 ------ 50 - x pounds
Mixture 2.72 ---------------- 50
Total 50 pounds
3.2 x + 2.4 ( 50 - x ) = 136
3.2 x + 120 - 2.4 x = 136
3.2 x - 2.4 x = 136 - -120
0.8 x = 16
/ 0.8
x = 20 pounds raisins
30 pounds nuts

m.ananth@hotmail.ca